# Structured Probabilistic Inference and Generative Modeling

### v2

@@ Abstract
Understanding causality in event sequences where outcome labels such as diseases or system failures arise from preceding events like symptoms or error codes is critical—yet remains an unsolved challenge across domains like healthcare or vehicle diagnostics. We introduce CARGO, a scalable multi-label causal discovery method for sparse, high-dimensional event sequences comprising of thousands of unique event types. Using two pretrained causal Transformers as domain-specific foundation models for event sequences—CARGO infers in parallel, per sequence one-shot causal graphs and aggregates them using an adaptive frequency fusion to reconstruct the global Markov boundaries of labels. This two-stage approach enables efficient probabilistic reasoning at scale while bypassing the intractable cost of full-dataset conditional independence testing. Our results on a challenging real-world automotive fault prediction dataset with over 29,100 unique event types and 474 imbalanced labels demonstrate CARGO’s ability to perform structured reasoning.
@@ 1 Introduction
Understanding why specific events lead to particular outcomes is vital for effective diagnosis, predictions and overall decision making [liu2025learning, shtp]. For instance, ”what serie of events captured by diagnostic led to this vehicle failure” or ”what symptoms led to this disease” [MANOCCHIO2024122564, MedBERT, bihealth, flight_service_cd, pdm_dtc_feature_extraction]. Here, an event sequence consists of a list of discrete events $x_{i}$ recorded asynchronously over time, while labels $\boldsymbol{y}$ summarize outcomes associated with the full sequence (e.g, a diagnosed defect or condition).
A fundamental obstacle in these settings is dimensionality. Real-world systems often involve tens of thousands of possible events, which renders the causal discovery intractable for current algorithms [hasan2023a]. To address this, we reinterpret multi-label causal discovery for event sequences as a form of Bayesian model averaging [HOET1999, pearl_1998_bn], where each sequence is treated as a sample from a local causal model. Specifically, each sequence induces a one-shot causal graph (i.e., a directed acyclic graph (DAG)) [tf_causalinterpretation_neurips_2023]. Together they can be fused to form a unified global structure [divide_conquer_mb_discovery]. This process, known as structural fusion [consensus_bn_jose], aggregates local graphs into a consensus causal graph over all observed sequences.
To summarise our contributions: we introduce CARGO (Causal Aggregation via Regressive Graph Operations), the first method to provide scalable causal discovery across thousands of labelled event sequences, with theoretical guarantees under standard assumptions. It is divided into two phases: (1) One-shot graph extraction, where for each sequence CARGO infers the local Markov boundary of each label using two autoregressive Transformers as density estimators [tf, density_estimator_2O21_journal_ci] (2) Graph fusion, where the local graphs are aggregated via an adaptive thresholding function to provide global Markov Boundaries.
We empirically validate CARGO on a large-scale vehicular dataset composed of about $29,100$ events and $474$ imbalanced labels, demonstrating for the first time, scalability and practical superiority over traditional causal discovery baselines. We also provide ablation on scoring criterions, frequency thresholds and the quality of the Transformers.
@@ 2 Preliminary & Related Work
Event Sequence Modelling. Event sequences are typically represented as a series of time-stamped discrete events $S=\{(t_{1},x_{1}),\ldots,(t_{L},x_{L})\}$ where $0\leq t_{1}<\ldots\leq t_{L}$ denotes the time of occurrence of event type $x_{i}\in\mathbb{X}$ drawn from a finite vocabulary $\mathbb{X}$. In multi-label settings, a binary label vector $\boldsymbol{y}\in\{0,1\}^{|\mathbb{Y}|}$ is attached to $S$ and indicates the presence of multiple outcome labels chosen from $\mathbb{Y}$ occuring at the final time step $t_{L}$. Together, this results in a multi-labeled sequence $S_{l}=(S,(\boldsymbol{y}_{L},t_{L}))$.
@@
Transformers [tf, gpt, touvron2023llamaopenefficientfoundation] have emerged as the dominant architecture for sequence modelling. Recent work leveraged Transformers in high-dimensional event spaces for next-event and label prediction. math2024harnessingeventsensorydata proposed a dual Transformer architecture where one model predicts the next event type (DTC), and the other predicts label occurrence (e.g, error pattern). Through this paper, we repurpose this dual architecture for causal discovery.
@@
Contemporary works point toward dividing classical causal discovery for high-dimensional datasets into sub-problems and graph aggregation. laborda_ring_based_distirbuted introduce a ring-based distributed algorithm for learning high-dimensional BN, divide_conquer_mb_discovery explores a distributed approach for large-scale causal structure learning and recursive_mb_approach_pmlr_large_scale for Markov Boundaries.
Bayesian Network [pearl_1998_bn] has served as a modelling technique for a variety of decision problems. It is defined as a triplet $\boldsymbol{U},\mathbb{G},P>$ with $P$ the joint distribution over a variable set $\boldsymbol{U}$ of a directed acyclic graph $\mathbb{G}=(\boldsymbol{U},E)$ with $E$ as the set of directed edges. This triplet must satisfy the Markov Condition: every random variable $U_{i}$ is independent of its non-descendant variables given its parents $\text{Pa}(U_{i})$ in $\mathbb{G}$. The directed edge $(U_{i}\rightarrow U_{j})$ encodes a probabilistic dependence. Thus, the joint probability distribution can be factorised as:
@@
Greedy Equivalence Search. (GES) [ges] is one of the most theoretically sound methods to recover a Markov equivalence class (MEC, Def. 4) of a DAG. Particularly in the context of large samples of data, GES provides theoretical guarantees of reaching the true graph. Formally, GES searches for the MEC of the graph $\mathbb{G}^{*}$ from the observational dataset $\boldsymbol{D}$ with distribution $p$. This defines the optimisation problem as:
@@
We will now explain the two phases of CARGO, which are (1) One-shot causal discovery (2) Graph aggregation using adaptive thresholding. The Proofs of Lemmas and Theorems can be found in the Appendix B.
@@ 3 One-Shot Causal Discovery
We would like to assess how much additional information event $X_{i}$ occurring at step $i$ provides about label $Y_{j}$ when we already know the past sequence of events $\boldsymbol{Z}=S_{<i}$. We essentially try to answer if:
@@
Density Estimation. We used two pretrained Transformers $(\text{Tf}_{x},\text{Tf}_{y})$ trained via maximum likelihood on a dataset of multi-labelled event sequences $D=\{S^{1}_{l},\cdots,S^{m}_{l}\}\subset\mathbb{S}$. We assume that they perfectly model the true conditional distributions of events and labels (A4). While due to the strict equivalences between CI-tests and conditional independence, it is difficult to provide a theoretical guarantee under imperfect models, we acknowledge that this assumption may be violated. We note that most of the causal discovery methods either have strong parametric data assumptions or assume a perfect CI-test. We provide an ablation study on the impact of the NADE’s quality on the one-shot phase in Appendix C.1, including number of parameters, context $c$ and $\text{Tf}_{y}$ performance.
@@
$\text{Tf}_{x}(S_{<i})$ $=\textit{Softmax}(\boldsymbol{h}^{x}_{i-1})=P_{\theta_{x}}(X_{i}|\boldsymbol{Z}),\;\hat{x}^{(z)}_{i}\triangleq P_{\theta_{x}}(X_{i}=x_{i}|\boldsymbol{Z})$ (7)
$\text{Tf}_{y}(S_{\leq i})$ $=\textit{Sigmoid}(\boldsymbol{h}^{y}_{i})=P_{\theta_{y}}(Y|X_{i},\boldsymbol{Z}),\hat{y}^{(z)}_{i+1,j}\triangleq P_{\theta_{y}}(Y=y_{j}|X_{i}=x_{i},\boldsymbol{Z})$ (8)
Sequential One-shot Causal Discovery. The CMI using Eq.(4) is computable only with the posteriors $P(Y_{j}|\boldsymbol{Z}),P(Y_{j}|X_{i},\boldsymbol{Z}))$. In practice a label-specific threshold $\theta_{j}\approx 0$ is applied to Eq (4) to identify conditional dependence:
@@
Computation. A key advantage of our approach is its scalability. Unlike traditional methods whose complexity depends on the event and label cardinality $|\mathbb{X}|$ and $|\mathbb{Y}|$ [feature_selection_review], phase 1 is agnostic to both. Figure 2 shows all parallelised steps on GPUs. CMI estimations are independently performed for all positions $i\in[c,L]$, with the sampling pushed into the batch dimension and results averaged across labels. This transitions the time complexity from $\mathcal{O}(\text{BS}\times N\times L)$ to $\mathcal{O}(1)$ per batch, with $L$ being the sequence length.
To ensure stable conditional entropy estimates and reliable predictions from $\text{Tf}_{y}$, the CMI is computed after observing $c$ events (context). This design choice also enables out-of-the-box parallelisation. By sampling $N$ variations of the prefix sequence $S_{\leq c}$, the CMI is independently computed across positions $i\in[c,L]$. In our experiments, we set $c=15,L=192$. The implementation of Phase 1 in Pytorch [pytorch] is provided in Appendix D.2.
Figure 2: The overview of CARGO. Phase 1 (One-shot) is on top, and Phase 2 (Adaptive Thresholding) is on the bottom. $d$ denotes the hidden dimension, $L$ the sequence length, $m$ the number of samples and $\text{MB}_{1},\text{MB}_{2}$ the Markov Boundary of $Y_{1},Y_{2}$. All green and blue areas are parallelised.
@@ 4 Structural Fusion of Markov Boundaries
Under the Oracle Models assumption (A4), our one-shot discovery phase acts as a perfect conditional independence tester. Consequently, the detection of an edge in a local graph $\mathbb{G}_{k}$ corresponds precisely to a true causal dependency in the global graph $\mathbb{G}^{*}$. The probability of this event $P(Z^{k}_{i,j}=1)$ is therefore the true marginal probability of the edge’s existence, which we denote as $\pi_{i,j}$.
@@ 4.1 Aggregation under imperfect CI-tests
The assumption of an Oracle CI-tester, while necessary for initial theoretical guarantees, is invariably violated in practice due to factors like model capacity, limited data or class imbalance. The extracted one-shot graphs will most likely violate the independencies in $\mathbb{G}_{k}$ and thus $\mathbb{G}^{*}$.
@@
The empirical frequency now converges to this new expectation. For a true edge $(\pi_{i,j}=1)$, the empirical frequency converges to a high value: $\hat{\pi}_{ij}(m)\xrightarrow{p}1-\beta$ and for spurious edge $(\pi_{i,j}=0)$ it converges to a low value: $\hat{\pi}_{ij}(m)\xrightarrow{p}\alpha$ This reveals the critical role of frequency aggregation as a mechanism for separating signal from noise.
@@ 4.2 Adaptive Fusion for Structural Discovery in Long-Tail Distributions
A primary challenge in real-world causal discovery is the long-tail distribution of outcome labels, where a few "head" labels possess abundant data while the vast majority of "tail" labels are data-sparse [longtailissue].
@@
This function smoothly interpolates between a user-defined maximum threshold, $\tau_{max}$ (prioritising precision for the tail), and a minimum, $\tau_{min}$ (prioritising recall for the head). Crucially, the function’s behaviour is calibrated by the data’s distribution. Such that decay midpoint $m_{0}$ is set to the median of all label supports, providing a robust anchor point against skew.
The decay rate, $k$ is made inversely proportional to the log-inter-quartile range of supports such that $k=\frac{2\log 3}{\log q_{75}-\log q_{25}}$. For rare labels with small $m_{j}$, the high variance of the frequency estimate necessitates a high threshold $\tau_{j}(m_{j})$ that acts as a strong regularizer. For common labels with large $m_{j}$, the LLN guarantees the convergence of $\hat{\pi}_{i,j}$ to the true edge probability, justifying a lower threshold to capture a more complete causal structure. Hence, this strategy effectively functions as a data-driven denoising mechanism and shares theoretical parallels with ensembling methods [beiman_decision_trees, consistency_rate_of_dc_tree] to enhance the robustness and accuracy of the final fused graph.
@@ 5 Empirical Evaluation
Settings & Vehicle Dataset. We used a $g4dn.12xlarge$ instance from AWS Sagemaker to run comparisons. It contains 48 vCPUs and 4 NVIDIA T4 GPUs. We used a combination of F1-Score, Precision, and Recall with different averaging [reviewmultilabellearning] to perform comparisons. We evaluated our method on a real-world vehicular test set of $m=300,000$ sequences, with $|\mathbb{Y}|=474$ different error patterns and $|\mathbb{X}|=29,100$ different DTCs forming sequences of $\approx 100\pm 35$ events. We used $\text{Tf}_{x}$ and $\text{Tf}_{y}$ with 90m and 15m parameters [math2024harnessingeventsensorydata]. The two NADEs didn’t see the test set during training. The error patterns are manually defined by domain experts as boolean rules between DTCs (Fig. 3). We set the elements of this rule as the correct Markov Boundary for each label $y_{j}$ in the tested sequences.
Multi-label Causal Discovery Comparisons. We benchmark CARGO against local structure learning (LSL) algorithms that estimate Markov Boundaries. This includes established approaches such as CMB [cmb], MB-by-MB [WANG2014252], PCD-by-PCD [pcdpcd], IAMB [iamb] from the PyCausalFS package [causality_based_feature_selection_2019], as well as the more recent, state-of-the-art MI-MCF [mimcf]. 6 random folds of the test data were created and converted into a multi-one-hot data-frame where one row represents one sequence and each column represents an event type or label $(\mathbb{X,Y})$.
Ablation on Aggregation Criterions for Phase 2. We provide an Ablation of the different criterion used in the structural fusion of Markov Boundaries. Union stands for a simple union over all edges without any removal. Frequency or edge voting, counts how often is $X_{i}\in\text{MB}(Y_{j})$. Then apply a static frequency threshold $\tau=[0.05,0.25,0.5,0.8]$. MI uses the mutual information between events and label as criterion in a Background Equivalence Search [ges]. Expected FPR (false positive ratio) [consensus_bn_freq_cutoff_no] describes two beta distributions which are fitted using the distribution of the mutual information $I(X_{i},Y_{j})$ extracted from Phase 1. The lower tail is used for outlier detection. Different FPR are chosen $\beta=[0.01,0.05,0.15,0.2]$. Detailed definitions can be found in Appendix C.4.

### v3

@@ Abstract
Understanding causality in event sequences where outcome labels such as diseases or system failures arise from preceding events like symptoms or error codes is critical—yet remains an unsolved challenge across domains like healthcare or vehicle diagnostics. We introduce CARGO, a scalable multi-label causal discovery method for sparse, high-dimensional event sequences consisting of thousands of unique event types. Using two pretrained causal Transformers as domain-specific foundation models for event sequences—CARGO infers in parallel, per sequence, one-shot causal graphs and aggregates them using an adaptive frequency fusion to reconstruct the global Markov boundaries of labels. This two-stage approach enables efficient probabilistic reasoning at scale while avoiding the intractable cost of full-dataset conditional-independence testing. Our results on a challenging real-world automotive fault prediction dataset with over 29,100 unique event types and 474 imbalanced labels demonstrate CARGO’s ability to perform structured reasoning.
@@ 1 Introduction
Understanding why specific events lead to particular outcomes is vital for effective diagnosis, predictions, and overall decision-making [liu2025learning, shtp]. For instance, ”what series of events captured by diagnostic led to this vehicle failure” or ”what symptoms led to this disease” [MANOCCHIO2024122564, MedBERT, bihealth, flight_service_cd, pdm_dtc_feature_extraction]. Here, an event sequence consists of a list of discrete events $x_{i}$ recorded asynchronously over time, while labels $\boldsymbol{y}$ summarize outcomes associated with the full sequence (e.g, a diagnosed defect or condition).
A fundamental obstacle in these settings is dimensionality. Real-world systems often involve tens of thousands of possible events, rendering causal discovery intractable for current algorithms [hasan2023a]. To address this, we reinterpret multi-label causal discovery for event sequences as a form of Bayesian model averaging [HOET1999, pearl_1998_bn], where each sequence is treated as a sample from a local causal model. Specifically, each sequence induces a one-shot causal graph (i.e., a directed acyclic graph (DAG)) [tf_causalinterpretation_neurips_2023]. Together, they can be fused to form a unified global structure [divide_conquer_mb_discovery]. This process, known as structural fusion [consensus_bn_jose], aggregates local graphs into a consensus causal graph over all observed sequences.
To summarize our contributions: we introduce CARGO (Causal Aggregation via Regressive Graph Operations), the first method to provide scalable causal discovery across thousands of labelled event sequences, with theoretical guarantees under standard assumptions. It is divided into two phases: (1) One-shot graph extraction, where for each sequence, CARGO infers the local Markov boundary of each label using two autoregressive Transformers as density estimators [tf, density_estimator_2O21_journal_ci] (2) Graph fusion, where the local graphs are aggregated via an adaptive threshold function to provide global Markov Boundaries.
We empirically validate CARGO on a large-scale vehicular dataset comprising about $29,100$ events and $474$ imbalanced labels, demonstrating, for the first time, scalability and practical superiority over traditional causal discovery baselines. We also perform ablation on scoring criteria, frequency thresholds, and Transformer quality.
@@ 2 Preliminary & Related Work
Event Sequence Modelling. Event sequences are typically represented as a series of time-stamped discrete events $S=\{(t_{1},x_{1}),\ldots,(t_{L},x_{L})\}$ where $0\leq t_{1}<\ldots\leq t_{L}$ denotes the time of occurrence of event type $x_{i}\in\mathbb{X}$ drawn from a finite vocabulary $\mathbb{X}$. In multi-label settings, a binary label vector $\boldsymbol{y}\in\{0,1\}^{|\mathbb{Y}|}$ is attached to $S$ and indicates the presence of multiple outcome labels chosen from $\mathbb{Y}$ occurring at the final time step $t_{L}$. Together, this results in a multi-labeled sequence $S_{l}=(S,(\boldsymbol{y}_{L},t_{L}))$.
@@
Transformers [tf, gpt, touvron2023llamaopenefficientfoundation] have emerged as the dominant architecture for sequence modelling. Recent work leveraged Transformers in high-dimensional event spaces for next-event and label prediction. math2024harnessingeventsensorydata proposed a dual Transformer architecture in which one model predicts the next event type (DTC) and the other predicts label occurrence (e.g., error patterns). Through this paper, we repurpose this dual architecture for causal discovery.
@@
Contemporary work points to decomposing classical causal discovery for high-dimensional datasets into sub-problems and graph aggregation. laborda_ring_based_distirbuted introduce a ring-based distributed algorithm for learning high-dimensional BN, divide_conquer_mb_discovery explores a distributed approach for large-scale causal structure learning and recursive_mb_approach_pmlr_large_scale for Markov Boundaries.
Bayesian Network [pearl_1998_bn] has served as a modelling technique for a variety of decision problems. It is defined as a triplet $\boldsymbol{U},\mathbb{G},P>$ with $P$ the joint distribution over a variable set $\boldsymbol{U}$ of a directed acyclic graph $\mathbb{G}=(\boldsymbol{U},E)$ with $E$ as the set of directed edges. This triplet must satisfy the Markov Condition: every random variable $U_{i}$ is independent of its non-descendant variables given its parents $\text{Pa}(U_{i})$ in $\mathbb{G}$. The directed edge $(U_{i}\rightarrow U_{j})$ encodes a probabilistic dependence. Thus, the joint probability distribution can be factorized as:
@@
Greedy Equivalence Search. (GES) [ges] is one of the most theoretically sound methods to recover a Markov equivalence class (MEC, Def. 4) of a DAG. In large-sample settings, GES provides theoretical guarantees of recovering the true graph. Formally, GES searches for the MEC of the graph $\mathbb{G}^{*}$ from the observational dataset $\boldsymbol{D}$ with distribution $p$. This defines the optimization problem as:
@@
We will now explain the two phases of CARGO, which are (1) One-shot causal discovery (2) Graph aggregation using adaptive thresholds. The Proofs of Lemmas and Theorems can be found in the Appendix B.
@@ 3 One-Shot Causal Discovery
We want to assess how much additional information event $X_{i}$ occurring at step $i$ provides about label $Y_{j}$ when we already know the past sequence of events $\boldsymbol{Z}=S_{<i}$. We essentially try to answer if:
@@
Density Estimation. We used two pretrained Transformers $(\text{Tf}_{x},\text{Tf}_{y})$ trained via maximum likelihood on a dataset of multi-labelled event sequences $D=\{S^{1}_{l},\cdots,S^{m}_{l}\}\subset\mathbb{S}$. We assume that they perfectly model the true conditional distributions of events and labels (A4). While, due to the strict equivalences between CI-tests and conditional independence, it is difficult to provide a theoretical guarantee under imperfect models, we acknowledge that this assumption may be violated. We note that most causal discovery methods either impose strong parametric data assumptions or rely on a perfect CI-test. We provide an ablation study on the impact of the NADE’s quality on the one-shot phase in Appendix C.1, including the number of parameters, context $c$, and $\text{Tf}_{y}$ performance.
@@
$\text{Tf}_{x}(S_{<i})$ $=\textit{Softmax}(\boldsymbol{h}^{x}_{i-1})=P_{\theta_{x}}(X_{i}|\boldsymbol{Z})$ (7)
$\text{Tf}_{y}(S_{\leq i})$ $=\textit{Sigmoid}(\boldsymbol{h}^{y}_{i})=P_{\theta_{y}}(Y|X_{i},\boldsymbol{Z})$ (8)
Sequential One-shot Causal Discovery. The CMI using Eq.(4) is computable only with the posteriors $P(Y_{j}|\boldsymbol{Z}),P(Y_{j}|X_{i},\boldsymbol{Z}))$. In practice a label-specific threshold $\theta_{j}\approx 0$ is applied to Eq. (4) to identify conditional dependence:
@@
Computation. A key advantage of our approach is its scalability. Unlike traditional methods whose complexity depends on the event and label cardinality $|\mathbb{X}|$ and $|\mathbb{Y}|$ [feature_selection_review], phase 1 is agnostic to both. Figure 2 shows all parallelized steps on GPUs. CMI estimations are independently performed for all positions $i\in[c,L]$, with the sampling pushed into the batch dimension and results averaged across labels. This transitions the time complexity from $\mathcal{O}(\text{BS}\times N\times L)$ to $\mathcal{O}(1)$ per batch, with $L$ being the sequence length.
To ensure stable conditional entropy estimates and reliable predictions from $\text{Tf}_{y}$, the CMI is computed after observing $c$ events (context). This design choice also enables out-of-the-box parallelization. By sampling $N$ variations of the prefix sequence $S_{\leq c}$, the CMI is independently computed across positions $i\in[c,L]$. In our experiments, we set $c=15,L=192$. The implementation of Phase 1 in PyTorch [pytorch] is provided in Appendix D.2.
Figure 2: The overview of CARGO. Phase 1 (One-shot) is on top, and Phase 2 (Adaptive Thresholding) is on the bottom. $d$ denotes the hidden dimension, $L$ the sequence length, $m$ the number of samples and $\text{MB}_{1},\text{MB}_{2}$ the Markov Boundary of $Y_{1},Y_{2}$. All green and blue areas are parallelized.
@@ 4 Structural Fusion of Markov Boundaries
Under the Oracle Models assumption (A4), our one-shot discovery phase serves as a perfect conditional-independence tester. Consequently, the detection of an edge in a local graph $\mathbb{G}_{k}$ corresponds precisely to a true causal dependency in the global graph $\mathbb{G}^{*}$. The probability of this event $P(Z^{k}_{i,j}=1)$ is therefore the true marginal probability of the edge’s existence, which we denote as $\pi_{i,j}$.
@@ 4.1 Aggregation under imperfect CI-tests
The assumption of an Oracle CI-tester, while necessary for initial theoretical guarantees, is invariably violated in practice due to factors like model capacity, limited data, or class imbalance. The extracted one-shot graphs will most likely violate the independencies in $\mathbb{G}_{k}$ and thus $\mathbb{G}^{*}$.
@@
The empirical frequency now converges to this new expectation. For a true edge $(\pi_{i,j}=1)$, the empirical frequency converges to a high value: $\hat{\pi}_{ij}(m)\xrightarrow{p}1-\beta$ and for a spurious edge $(\pi_{i,j}=0)$ it converges to a low value: $\hat{\pi}_{ij}(m)\xrightarrow{p}\alpha$ This reveals the critical role of frequency aggregation as a mechanism for separating signal from noise.
@@ 4.2 Adaptive Fusion for Structural Discovery in Long-Tail Distributions
A primary challenge in real-world causal discovery is the long-tail distribution of outcome labels, where a few “head” labels possess abundant data while the vast majority of “tail” labels are data-sparse [longtailissue].
@@
This function smoothly interpolates between a user-defined maximum threshold, $\tau_{max}$ (prioritizing precision for the tail), and a minimum, $\tau_{min}$ (prioritizing recall for the head). Crucially, the function’s behavior is calibrated by the distribution of the data. Such that decay midpoint $m_{0}$ is set to the median of all label supports, providing a robust anchor point against skew.
The decay rate, $k$ is made inversely proportional to the log-inter-quartile range of supports such that $k=\frac{2\log 3}{\log q_{75}-\log q_{25}}$. For rare labels with small $m_{j}$, the high variance of the frequency estimate necessitates a high threshold $\tau_{j}(m_{j})$ that acts as a strong regularizer. For common labels with large $m_{j}$, the LLN guarantees the convergence of $\hat{\pi}_{i,j}$ to the true edge probability, justifying a lower threshold to capture a more complete causal structure. Hence, this strategy serves as a data-driven denoising mechanism and shares theoretical parallels with ensembling methods [beiman_decision_trees, consistency_rate_of_dc_tree], thereby enhancing the robustness and accuracy of the final fused graph.
@@ 5 Empirical Evaluation
Settings & Vehicle Dataset. We used a $g4dn.12xlarge$ instance from AWS Sagemaker to run comparisons. It contains 48 vCPUs and 4 NVIDIA T4 GPUs. We used a combination of F1-Score, Precision, and Recall with different averaging methods [reviewmultilabellearning] to compare results. We evaluated our method on a real-world vehicular test set of $m=300,000$ sequences, with $|\mathbb{Y}|=474$ different error patterns and $|\mathbb{X}|=29,100$ different DTCs forming sequences of $\approx 100\pm 35$ events. We used $\text{Tf}_{x}$ and $\text{Tf}_{y}$ with 90m and 15m parameters [math2024harnessingeventsensorydata]. The two NADEs didn’t see the test set during training. Domain experts manually define error patterns as Boolean rules among DTCs (Fig. 3). We set the elements of this rule as the correct Markov Boundary for each label $y_{j}$ in the tested sequences.
Multi-label Causal Discovery Comparisons. We benchmark CARGO against local structure learning (LSL) algorithms that estimate Markov Boundaries. This includes established approaches such as CMB [cmb], MB-by-MB [WANG2014252], PCD-by-PCD [pcdpcd], IAMB [iamb] from the PyCausalFS package [causality_based_feature_selection_2019], as well as the more recent, state-of-the-art MI-MCF [mimcf]. 6 random folds of the test data were created and converted into a multi-one-hot data-frame where one row represents one sequence, and each column represents an event type or label $(\mathbb{X,Y})$.
Ablation on Aggregation Criterions for Phase 2. We provide an Ablation of the different criteria used in the structural fusion of Markov Boundaries. Union stands for a simple union over all edges without any removal. Frequency or edge voting, counts how often is $X_{i}\in\text{MB}(Y_{j})$. Then apply a static frequency threshold $\tau=[0.05,0.25,0.5,0.8]$. MI uses the mutual information between events and labels as a criterion in a Background Equivalence Search [ges]. Expected FPR (false positive ratio) [consensus_bn_freq_cutoff_no] describes two beta distributions which are fitted using the distribution of the mutual information $I(X_{i},Y_{j})$ extracted from Phase 1. The lower tail is used for outlier detection. Different FPR are chosen $\beta=[0.01,0.05,0.15,0.2]$. Detailed definitions can be found in Appendix C.4.